OpenStudy (anonymous):

how to make (5,6),(2,0),(0,-4),and(-2,-8) into an equation

OpenStudy (anonymous):

is it geometry???

OpenStudy (anonymous):

some details regarding question please.

OpenStudy (anonymous):

if it is a line you can

OpenStudy (anonymous):

equation of straight line?

OpenStudy (anonymous):

i can. but can you post screenshot of the question?

OpenStudy (anonymous):

is it one equation or two??

OpenStudy (anonymous):

do u know about slope?

OpenStudy (phi):

Do you know how to find the equation of a line, when you are given two points?

OpenStudy (anonymous):

yes y=ax+b and a=f(x1)-f(x2)/x1-x2

OpenStudy (anonymous):

It's a linear relation; only two points are needed. Start with slope formula, \(\large m=\frac{y-y_1}{x-x_1}\) Then you can rearrange that to \(\large y-y_1=m(x-x_1)\)